OpenStudy (anonymous):

Write an algebraic expression that describes this set of calculations. step 1- take some number and square it. step 2- add 2 times the starting number step 3- divide that result by your starting number. step 4- subtract your starting number.

OpenStudy (jamesj):

What the equation for step 1?

OpenStudy (jamesj):

call the number x

OpenStudy (anonymous):

i dont know. it just says that we need to take any number.

OpenStudy (jamesj):

Yes, so call that arbitrary starting number x. The first step is x^2

OpenStudy (jamesj):

What now is the second step?

OpenStudy (anonymous):

add two times the starting number

OpenStudy (jamesj):

yes, in terms of x ...

OpenStudy (anonymous):

huh?

OpenStudy (jamesj):

x is the starting number. The first step is that number squared. That is: \[ x^2 \]

OpenStudy (jamesj):

Using x, the starting number, what now is the next step?

OpenStudy (anonymous):

2(x^2)?

OpenStudy (jamesj):

No. You need to add 2 times the starting number. What is just 2 times the starting number, in terms of x?

OpenStudy (anonymous):

you dont make sense!

OpenStudy (jamesj):

It's \[ 2x \]

OpenStudy (anonymous):

for step one or step two

OpenStudy (jamesj):

We have x^2 from step one. And step two says now add 2 times the starting number, i.e., 2x. So what do we have as an algebraic expression after step 2. That is, step 1 and 2.

OpenStudy (anonymous):

so for step 2 i need to write x^2+ 2x

OpenStudy (jamesj):

Yes. Now step 3?

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

How do you do it? I'm asking you. I know, I want to see if you get the idea now.

OpenStudy (anonymous):

so next is going to be x^2+2x over x ?

OpenStudy (jamesj):

yes

OpenStudy (jamesj):

and that expression simplifies to what?

OpenStudy (anonymous):

x^1+2

OpenStudy (jamesj):

Yes, or just x + 2. Ok, now step 4....

myininaya (myininaya):

(a different problem) Example: Step 1: Take some number and take the square root of it: (let's call this number x) so we need the square root of x: \[\sqrt{x}\] Step 2: Subtract a half of the starting number \[- \frac{1}{2} \cdot x\] Step 1 together with Step 2: so far we have \[\sqrt{x}-\frac{1}{2}x\] Step 3: Multiply the result by your starting number: \[x(\sqrt{x}-\frac{1}{2}x)\] Step 4: add your starting number \[+x\] Step 3 together with Step 4: \[x(\sqrt{x}-\frac{1}{2}x)+x\]

OpenStudy (anonymous):

before we do step 4 , how do i know i need to simplify

OpenStudy (jamesj):

Because you always should if you can.

OpenStudy (jamesj):

It's a general rule in mathematics

OpenStudy (jamesj):

But if you're nervous about it, don't.

OpenStudy (anonymous):

this just made me feel so dumb :D lol. okay step 4: x+2-x

OpenStudy (jamesj):

and that simplifies to .... what?

OpenStudy (anonymous):

2?

OpenStudy (jamesj):

Yes. Just 2.

OpenStudy (anonymous):

all this work just for the number two -_- oh lord.

OpenStudy (jamesj):

So in summary: \[{x^2 + 2x \over x} - x \ = \ 2\]

myininaya (myininaya):

and that will =2 for any number except for x=0

OpenStudy (jamesj):

The object of the question obviously isn't to write down "2", but to show you how these things are done. And the fact that it does reduce to 2 is actually the teacher/book's way of telling you you've got it right, because the final answer is so simple. Sadly this isn't the case in real life, but it's great that it is when you're learning ;-)

OpenStudy (anonymous):

true. thanksss:D